$6,000 dollars is invested in two different accounts earning 3% and 5% interest. At the end of one year, the two accounts earned $220 in interest. How much money was invested at 5%?
Answer by mananth(16949)


Part I 5.00% per annum ------------- Amount invested =x
Part II 3.00% per annum ------------ Amount invested = y
6000
Interest----- 220.00

Part I 5.00% per annum ---x
Part II 3.00% per annum ---y
Total investment
x + 1 y= 6000 -------------1
Interest on both investments
5.00% x + 3.00% y= 220
Multiply by 100
5 x + 3 y= 22000.00 --------2
Multiply (1) by -5
we get
-5 x -5 y= -30000.00
Add this to (2)
0 x -2 y= -8000
divide by -2
y = 4000
Part I 5.00% $ 2000
Part II 3.00% $ 4000

CHECK
2000 --------- 5.00% ------- 100.00
4000 ------------- 3.00% ------- 120.00
Total -------------------- 220.00

m.ananth@hotmail.ca